OpenStudy (anonymous):

A bag contains 6 red balls and 5 green balls. You randomly draw one ball, replace it, then randomly draw another ball. How would you solve this?

OpenStudy (agreene):

6+5 = 11 (total balls) 6/11 = likelihood of getting a red ball 5/11 = likelihood of getting a blue ball 6/11 = 0.5454 or 54.54% chance of getting red ball 5/11 = 0.4545 or 45.45% chance of getting blue ball.